Water is pouring into a cuboidal reservoir at the rate of 60 liters per minute. If the volume of reservoir is 108 m^3, find the number of hours it will take to fill the reservoir.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given, volume of reservoir = 108m^3
Rate of pouring water into cuboidal reservoir = 60 liters/minute
= 60/1000m^3 per minute
Since 1 liter = (1/1000 )m^3
= (60×60)/1000m^3per hour
Therefore, (60×60)/1000m^3water filled in reservoir will take = 1 hour
Therefore1m^3water filled in reservoir will take = 1000/(60×60) hours
Therefore,108m^3water filled in reservoir will take = (108×1000)/(60×60) hours = 30 hours
Answer: It will take 30 hours to fill the reservoir.

Thanks